OpenStudy (anonymous):

find the average value of the function over the given interval. f(x)=4(x^2+1)/x^2, [1,3]

OpenStudy (anonymous):

\[\frac{ 1 }{ 3-1 }*\int\limits_{1}^{3}\frac{ 4*(x^2+1) }{ x^2 }\]

OpenStudy (anonymous):

dx

OpenStudy (anonymous):

what does the \[\frac{ 4(x ^{2}+1) }{x ^{2} ? }\] integrate into?

OpenStudy (anonymous):

727 269 1800

OpenStudy (anonymous):

?